A refrigerator works on the carnet cycle in temperature between -7°C and 27'C. It makes 500 kg of ice per hour at-5'C from water at 14'C. Find H.P required to drive the compressor and C.O.P. of the cycle. Take specific heat of ice as 2.1 KJ/Kg-k and latent heat as 336 kJ/Kg


1
Expert's answer
2021-05-17T17:05:01-0400

COP = "\dfrac{Q}{W} = \dfrac{T_C}{T_H-T_C}"


TH = 27°C = 273K + 27 = 300K

TC = -7°C = 273K - 7 = 266k


COP = "\dfrac{266}{300-266}" = "\dfrac{266}{34}" = "7.824"